OpenStudy (anonymous):

Please evaluate the integral: (sin^2)9xdx

OpenStudy (anonymous):

do you mean \[\int\limits_{}^{}\sin^2(9x)dx\]?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

I know you have to use a trig indentity, sin^2= 1/2 (1-cos^2) but after that I'm alittle shakey

OpenStudy (anonymous):

that is the right way to start, although it's sin^2(x) = (1-cos^2(2x))/2, if we apply that to our integral, we get: \[(1/2)\int\limits_{}^{}(1- \cos^2(18x))\]. The integral of 1 is simple, but be sure to watch out for the 18x inside the cosine: \[= (x - [\sin(18x)/18])/2 + C\] (don't forget +C )

OpenStudy (anonymous):

Thanks for your help, I really do appreciate it. : ) I see where I went wrong, and I didn't think about taking the 1/2 out.

OpenStudy (anonymous):

Correction here. It's 1/2(1 - cos(2x)), without the square.

OpenStudy (anonymous):

^ thanks. He's right. The identity is without the square on the cosine. pretty sure everything else is good?

OpenStudy (anonymous):

I believe it is.

OpenStudy (anonymous):

Opps, that was also my bad. Sorry about that.